# Using Biot-Savart boundary conditions for unbounded external flow on Eulerian meshes

### v2

@@ Abstract
We introduce a novel boundary condition formulated using a Biot-Savart vorticity integral inside an Eulerian fluid domain that maintains high-accuracy results even when the domain boundary is within a body-length of immersed solid boundaries. The Biot-Savart condition is incorporated into a general solver with Eulerian velocity-pressure fields by including the pressure-induced body vorticity during the incompressible-flow projection-step. We use a multilevel approach to reduce the computational cost of the evaluation of the Biot-Savart integral from $O(N)$ to $O(\log N)$ and formally show that the error associated with this approach is bounded and scales with the inverse power of the problem’s dimension. We use two-dimensional and three-dimensional analytic vortex problems to demonstrate this scaling is correct and select appropriate multilevel coarsening parameters accordingly. We show that the new method exactly captures the analytical added-mass force of accelerating plates and matches experimentally measured wake development even when the entire domain only extends 1/2 diameter from the plate. The new method also predicts essentially identical time-varying drag forces for a two-dimensional circle regardless of domain size, while classical boundary conditions require a domain nearly 48 times larger to converge on the new method’s result. Finally, we study the sensitivity of 2D deflected wakes to the new domain boundary conditions and show that only in minimal domains, 25 times smaller than the reference domain, the omission of the far-field wake influences the flow field and the forces on the body. Doubling the domain size recovers the correct results.
@@
[inst1]organization=Ship Hydromechanics,addressline=Faculty of Mechanical Engineering (ME),
@@
Unbounded flow simulations on minimal Eulerian domains via Biot-Savart boundary conditions.
Fast ($\log N$) multilevel evaluation of the Biot-Savart integral with controllable bounded error.
Fully-couples velocity-pressure description to vorticity-condition via projection step.
Matches analytic and experimental results while reducing required domain size by 48x.
@@ 1 Introduction
Computational fluid dynamics provides a numerical solution to a continuum problem in an (often) unbounded domain. Common practice is to approximate the continuum problem with a discrete set of equations and truncate the unbounded domain to a finite computational domain with far-field boundary conditions.
External incompressible flows enforce far-field boundary conditions on the computational domain’s exterior. These consist of a known velocity normal to the boundary and a known normal pressure gradient111for a description of the boundary conditions for the incompressible Navier-Stokes equations, we refer the reader to [1]. These can lead to significant errors. In particular, the force acting on a body can suffer from severe blockage effects when the computational domain is not sufficiently large. Here, sufficiently large means that the decaying potential flow induced by the body and the advection/diffusion of vorticity through the domain’s boundary have no influence on the solution [2]. While domain sensitivity study can determine the extent of the blockage, there is no a priori rule for estimating their magnitude, making this an expansive iterative process. It is also computationally inefficient to add grid cells to fill such a large domain, and grid refinement near the solid boundaries introduces well-known sensitives to the grid expansion ratios and cell aspect ratios due to numerical diffusion, dispersion, reflection, and linear system stiffness. Alternatively, numerical dissipation can be explicitly imposed in a buffer layer on a smaller domain that forces the solution to decay rapidly before it reaches the boundaries [3].
Domain augmentation is a more principled approach to simulate external flow, such as using a large coarse grid to provide artificial far-field boundary conditions for a small well-resolved domain [2]. The outer flow field can also be augmented with a vortex particle-mesh method, reducing dispersive errors [4]. Such methods enable tracking the wake for an extended time while enabling accurate modeling of the developing boundary layer on the object’s surface and can also include immersed boundaries [5]. However, several (user-defined) parameters are required to correctly interpolate between the two domains, and the complexity of accurately coupling the domains makes theis apporach difficult to apply to existing flow solvers.
For flows with one [6] or two unbounded directions [7, 8], coordinate transforms and/or analytical potential flow solution [9] can be applied in the unbounded directions. Potential flow boundary conditions are prescribed following the classical Helmholtz decomposition of the velocity field into an irrotational and an incompressible part
@@
where $\vec{u}_{\omega}$, $\nabla\phi$, $\vec{U}_{\infty}$ are the incompressible component, irrotational component, and the free-stream velocity, respectively. While these methods are elegant, they rely either on a single unbounded direction (as in [6] or formulating the problem in a coordinate system where two unbounded directions are transformed to a single one (for example, from Cartesian to Cylindrical in [8]), rendering them difficult to generalize to 3D flow problems.
@@
Other classes of methods solve unbounded Poisson equations by incorporating the unbounded conditions as Dirichlet boundary data, either by successive relaxed iterations of the Poisson problem with the Dirichlet data computed using the previous solution [10] or by using the James-Lackner [11, 12] algorithm that solves two nested domain Poisson problems, the inner-most with homogeneous Neumann boundary conditions and second outermost domain with inhomogeneous Neumann boundary conditions computed from a boundary convolution of the inner solution with the fundamental solution to the Poisson operator. To accelerate the convolution, multilevel approximation [13] or using local correction methods [14] can be used. Although these methods enforce unbounded boundary conditions, they require iterating a Poisson problem or solving it multiple times.
@@
Finally, fundamental solutions to discrete operators (Lattice Green’s functions or LGFs) and their asymptotic expansions can be used to solve the discrete Poisson equations in unbounded domains [15, 16]. Points close to the source are evaluated with the discrete kernel (accurate), while points far away are evaluated with the asymptotic expansion (fast). These methods rely on the unbounded properties of the LGFs and can simulate unbounded viscous incompressible flow [17] and flows with immersed-boundaries [18]. This allows for computational domains snug to the vorticity field, but adapting these methods to to existing Eulerian or immersed boundary solvers is difficult.
@@
In this manuscript, we propose a novel method for imposing external potential-flow boundary conditions of a viscous flow simulation by coupling to a vorticity-based description of the flow on the boundary via the Biot-Savart equation. The method retains the primitive variables formulation ($\vec{u}$ and $p$) of the governing equations, is compatible with immersed boundary methods and fast pressure projection methods, and does not require additional meshes to describe the flow. In the next section, we describe the Biot-Savart boundary conditions and how they are injected into a projection algorithm to solve the Navier-Stokes equations. Next, we demonstrate the accuracy of the velocity reconstruction and the balance between speedup and accuracy in the reconstructed field. With confined vorticity and vortex wake examples, we demonstrate that the new boundary conditions significantly outperform standard reflective boundary conditions on significantly smaller domains. We present an unexpected result: the new method continues to perform well when the fundamental assumption of potential flow external to the domain is violated, and we use the unstable deflected wakes generated by fast flapping foils to study the limits of the approach. We finish this manuscript by discussing the method’s applicability to various flows in science and engineering.
@@ 2 Method
Figure 1: Schematic of the fluid domain $\Omega$ with an immersed body $\mathcal{B}$ and their common (wet) interface $\partial\mathcal{B}$. The exterior of the computational domain $\Omega$ is denoted as $\partial\Omega$ with $\hat{n}$ the unit normal vector.
We start by describing the governing equations and their numerical solution. The flow is governed by the Navier-Stokes equations for an incompressible flow. We discretize these equations on a staggered finite-volume mesh. The resulting system of equations is solved within a computational domain $\Omega(N)$, where $N$ is the number of degrees of freedom of the problem. The outer domain boundary is denoted at $\partial\Omega(N^{\cal S})$, where $\cal S$ is a factor linked to the dimensions of the problem. In 2D we have ${\cal S}=1/2$, and in 3D ${\cal S}=2/3$. The primary source of vorticity in the domain is an immersed body $\mathcal{B}$ with outer (wet) surface $\partial\mathcal{B}$ and unit normal vector $\hat{n}$, see Figure 1.
@@
We use the standard projection scheme [19] to solve the coupled velocity-pressure system inside a predictor-correct method to achieve second-order temporal convergence of the flow variables [20]. To simplify the discussion, we present the method for the first step of the predictor-corrector; generalization to the second-order Heun’s corrector method or higher-order time integration method is straightforward. The projection method starts with an explicit estimate of the intermediate velocity field $u^{*}$
@@
where Re is the Reynolds number of the flow. Depending on the choice of integration method for the right-hand-side, we obtain explicit or semi-implicit methods. Here we will focus on fully explicit methods. On the immersed body $\cal B$, the intermediate velocity is subject to the condition
$\vec{u}^{*}=\vec{U}_{b}\quad\forall x\in\partial\cal{B}.$ (3)
where $\vec{U}_{b}$ is the body velocity. Body-fitted methods enforce this condition directly, while immersed boundary methods enforce it indirectly; either by reformulating the governing equations near the boundary such as [21, 22] or by modifying the discretization near $\cal{B}$ [23].
Additionally, Eq. (2) must be supplemented with far-field boundary conditions on the domain’s outer boundaries. The simplest of these is a reflection condition
$\frac{\partial\vec{u}^{*}_{s}}{\partial\hat{n}}=0,\ \vec{u}^{*}_{n}=U_{n}$ $\quad\forall\ \vec{x}\in\partial\Omega(N^{\cal S}),$ (4)
where we have decomposed the velocity field on the boundary in a boundary normal $\vec{u}^{*}_{n}$ and tangential part $\vec{u}^{*}_{s}$. We refer the reader to [1] for a discussion of pressure boundary conditions on the domain’s exterior.
@@
As mentioned previously, the boundary $\partial\Omega$ must be placed sufficiently far away from the body not to influence the fast decaying (potential flow) part of the velocity field and ensure enough space to properly advect the vorticity downstream.
2.1 Biot-Savart boundary conditions
We propose to replace the typical “local” domain boundary conditions such as Eq. (4) &  (7) by a Biot-Savart integral over the vorticity inside the domain. A Green’s function solution to the vorticity equation Eq. (1) gives the velocity at a point $\vec{x}$ as
${\vec{u}}({\vec{x}})=f(\vec{x};\vec{\omega},\Omega)=\int_{\Omega}K_{n}({\vec{x}}-\vec{y})\times\vec{\omega}({\vec{y}})\text{ d}\vec{y}\quad\quad\forall\vec{x}\,\in\partial\Omega(N^{\cal S})$ (8)
This formulation implicitly assumes the flow external to the domain is irrotational and that the flow is incompressible everywhere. This assumption is met when all the vorticity is confined within the computational domain such as for the flow around accelerated bodies when the time $\frac{tU}{L}\lesssim t^{\text{crit.}}$, with $t^{\text{crit.}}$ the critical time for vortex detachment [25]. However, we will show that the results of using equation (8) in an Eulerian solver are still highly accurate for flows with wakes that extend to the domain boundary.
Note that on a staggered grid, we only need to apply the integral equation (8) to compute the velocity normal to the domain face. For ghost cells external to the computational domain, we use the local derivative conditions $\nabla\cdot\vec{u}=\nabla\times\vec{u}=0$ to compute the velocity. These conditions are consistent with the assumption that the flow is potential outside the domain and fast to apply. Specifically, we apply the zero curl condition to all tangential velocity components in the ghost cells and then the zero divergence condition to the normal components in the ghost cells, marching away from the domain boundary.

### v3

@@ Abstract
[…]
@@
[inst1]addressline=Faculty of Mechanical Engineering (ME),
@@
Biot-Savart boundary conditions produce accurate unbounded flow simulations on Eulerian domains minimally larger than the immersed solid geometries.
Iterative partitioned solution of the non-locally coupled immersed surfaces and domain surfaces in the projection step requires minimal adjustment to classic velocity-pressure incompressible Navier-Stokes solvers.
Simplified oct-tree clustering for the Fast Multi-level Method ($\ell$M) integral evaluation has bounded error, optimal $O(N)$ operations, and efficient parallel execution on GPUs.
@@ 1 Introduction
When simulating external viscous flow problems with an Eulerian method using velocity and pressure as the primary variables, the ideally unbounded domain must be truncated to a finite size and artificial domain boundary conditions must be imposed on the computational domain’s exterior. These typically consist of a known velocity normal to the boundary and a known normal pressure gradient [1]. These can lead to significant blockage errors when the computational domain is not sufficiently large, restricting the decay of the potential flow induced by the body, or advecting/diffusing vorticity through the domain’s boundary [2]. While domain sensitivity study can determine the extent of the blockage, there is no a priori rule for estimating their magnitude, and it is computationally inefficient to add grid cells to fill a large domain.
To reduce blockage effects and computational cost, domain augmentation method use a large coarse grid to provide artificial boundary conditions to a small well-resolved domain [2]. The outer flow field can also be augmented with a vortex particle-mesh method, reducing dispersive errors [3]. Such methods enable tracking the wake for an extended time while enabling accurate modeling of the developing boundary layer on the object’s surface and can also include immersed boundaries [4]. However, these methods rely on a coupling of an external Lagrangian solver to an internal Eulerian solver formulated in different variables.
An alternative approach is to utilize some form of potential flow boundary conditions, following the classical Helmholtz decomposition of the velocity field $\vec{u}(\vec{x})$ into an irrotational and an incompressible part
@@
where $\vec{u}_{\omega}$, $\nabla\phi$, $\vec{U}_{\infty}$ are the incompressible, irrotational, and the free-stream components, respectively. While these methods are elegant, they rely either on a single unbounded direction as in [5] or formulating the problem in a coordinate system where two unbounded directions are transformed to a single one (for example, from Cartesian to Cylindrical in [6]), which do not occur regularly in practical flow applications.
@@
A related class of methods solves unbounded Poisson equations by incorporating the unbounded conditions as Dirichlet boundary data, either by successive relaxed iterations of the Poisson problem with the Dirichlet data computed using the previous solution [7] or by using the James-Lackner [8, 9] algorithm that solves two nested domain Poisson problems, the inner domain with homogeneous Neumann boundary conditions and outer domain with inhomogeneous Neumann boundary conditions computed from a boundary convolution of the inner solution with the fundamental solution to the Poisson operator. To accelerate the convolution, multilevel approximation [10] or local correction methods [11] can be used.
@@
Finally, fundamental solutions to discrete operators (Lattice Green’s functions or LGFs) and their asymptotic expansions can be used to solve the discrete Poisson equations on unbounded domains [12, 13]. These methods rely on the unbounded properties of the LGFs and can simulate unbounded viscous incompressible flow [14] and flows with immersed-boundaries [15]. This allows for snug computational domains limited to regions of non-negligible vorticity or possibly a set of disjoint domains limited to active vorticity regions. Combined with a multi-resolution approach, the LGF approach can yield fast and accurate computation of 3D incompressible flows at a cost reduced compared to Cartesian meshes [16]. However, the method relies on a structured grid and a constant coefficient Poisson equation, and the possibility of extending the method to unstructured mesh and variable coefficient Poisson equation problems (such as the one arising in [17]) is unclear.
@@
In this manuscript, we propose a novel method for imposing unbounded external flow boundary conditions on a finite Eulerian domain via a Biot-Savart integral solution applied on the domain’s boundary. The method retains the primitive variables formulation ($\vec{u}$ and the pressure $p$) of the governing equations, is compatible with immersed boundary methods, variable coefficient Poisson equations, fast pressure projection methods, and does not require additional meshes to describe the flow. In the next section, we describe the Biot-Savart boundary conditions and how they are injected into a projection algorithm to solve the Navier-Stokes equations. Next, we demonstrate the insensitivity of this unique approach to errors in the velocity reconstruction on the boundary, and use that to develop a simplified multi-level clustering approach for an optimal $O(N)$ Fast Multi-level Method to evaluate the Biot-Savart integral. We validate the correctness of the coupled equations using accelerating flow examples where the external flow is truly irrotational and demonstrate that the new boundary conditions vastly outperform standard reflective boundary conditions on significantly smaller domains. We present an unexpected result: the new method continues to perform well when the fundamental assumption of potential flow external to the domain is violated, and we use the unstable deflected wakes generated by fast flapping foils to study the limits of the approach. We finish this manuscript by discussing the method’s applicability to various flows in science and engineering.
@@ 2 Method
Figure 1: Schematic of the fluid domain $\Omega$ with an immersed body $\mathcal{B}$ and their common (wet) interface $\partial\mathcal{B}$. The computational domain $\Omega$ has external boundary $\partial\Omega$ with $\hat{n}$ the unit normal vector.
The flow is governed by the incompressible Navier-Stokes equations
$\begin{array}[]{c}\rho\frac{\partial\vec{u}}{\partial t}+\rho\left(\vec{u}\cdot\nabla\right)\vec{u}=-\nabla p+\mu\nabla^{2}\vec{u}\\ \nabla\cdot\vec{u}=0\end{array}$ $\quad\quad\forall\ \vec{x}\in\Omega,$ (4)
where $\rho,\mu$ are the fluid’s density and viscosity and $\Omega$ is the fluid domain , see Fig. 1. These equations are supplemented with the no-slip boundary condition on the immersed body $\cal B$ with boundary $\partial\cal{B}$
$\vec{u}=\vec{U}_{b}\quad\forall x\in\partial\cal{B}$ (5)
and a velocity condition on the domain boundary $\partial\Omega$
$\vec{u}(\vec{x})=f(\vec{x},\vec{U}_{\infty})$ $\quad\forall\ \vec{x}\in\partial\Omega,$ (6)
where $\vec{U}_{b}$ is the body velocity and $f$ is the imposed condition and $\vec{U}_{\infty}$ is the far-field velocity. This work focuses on simulating external flow, i.e. $f$ is the identity function applied on an infinite domain $\partial\Omega\rightarrow\infty$.
There are many possible conditions to apply on a finite domain to simulate the external conditions, and the simplest of these is a (local) reflection condition
$\frac{\partial\vec{u}_{s}}{\partial\hat{n}}=0,\ \vec{u}_{n}=\vec{U}_{\infty}\cdot\hat{n}$ $\quad\forall\ \vec{x}\in\partial\Omega,$ (7)
where we have decomposed the velocity field on the boundary in a boundary normal $\vec{u}_{n}$ and tangential part $\vec{u}_{s}$. We refer the reader to [1] for a discussion of pressure boundary conditions on the domain’s exterior. As mentioned previously, the boundary $\partial\Omega$ must be placed sufficiently far away from the body not to influence the decaying potential flow part of the velocity field and ensure enough space to properly advect the vorticity downstream.
We propose to replace the typical “local” domain boundary conditions such as Eq. 7 by a Biot-Savart integral over the vorticity inside the domain. A Green’s function solution to the vorticity equation Eq. 1 gives the velocity at a point $\vec{x}$ as
${\vec{u}}({\vec{x}})=\vec{U}_{\infty}+f(\vec{x};\vec{\omega},\Omega)=\vec{U}_{\infty}+\int_{\Omega}K_{n}({\vec{x}}-\vec{y})\times\vec{\omega}({\vec{y}})\text{ d}\vec{y}\quad\quad\forall\vec{x}\,\in\partial\Omega$ (8)
The vortex-induced velocity of the Biot-Savart integral is simply added to the uniform background velocity $\vec{U}_{\infty}$ to determine the total velocity at any point. We will omit further discussion of the trivial uniform flow contribution in this section.
This formulation implicitly assumes the flow external to the domain is irrotational and that the flow is incompressible everywhere. This assumption is met when all the vorticity is confined within the computational domain such as for the flow around accelerated bodies when the time $\frac{tU}{L}\lesssim t^{\text{crit.}}$, with $t^{\text{crit.}}$ the critical time for vortex detachment [19]. However, we will show that the results of using Eq. 8 in an Eulerian solver are still highly accurate for flows with wakes that extend to the domain boundary.
Note that on a staggered grid, we only need to apply the integral equation (Eq. 8) to compute the velocity normal to the domain face. For ghost cells external to the computational domain, we use the local derivative conditions $\nabla\cdot\vec{u}=\nabla\times\vec{u}=0$ which are consistent with the assumption that the flow is potential outside the domain. Specifically, once we have used Eq. 8 to set $u_{n}$ along the domain boundary we apply the zero curl condition to all tangential velocity components in the ghost cells and then the zero divergence condition to fill the remaining normal component each ghost cell. For example, in 2D we sequentially apply
$\frac{\partial\vec{u}_{s}}{\partial\hat{n}}=\frac{\partial\vec{u}_{n}}{\partial\hat{s}}\quad\text{and}\quad\frac{\partial\vec{u}_{n}}{\partial\hat{n}}=-\frac{\partial\vec{u}_{s}}{\partial\hat{s}}.$
This process can continue for multiple layers of ghost cells, marching away from the domain boundary.
There are two primary issues with the use of Eq. 8 as the domain boundary condition for the incompressible Eulerian solver: (i) projecting the velocity into a divergence-free field consistent with the Biot-Savart equation, and (ii) the fast evaluation of Eq. 8. We demonstrate in the following section that these issues may be efficiently and accurately overcome largely without modifying the method used to solve the Navier-Stokes equation.
2.1 Biot-Savart Projection Algorithm
@@
We use the standard projection scheme [20] to solve the coupled velocity-pressure system inside a predictor-correct method to achieve second-order temporal convergence of the flow variables [17]. To simplify the discussion, we present the method for the first step of the predictor-corrector; generalization to the second-order Heun’s corrector or higher-order time integration method is straightforward. The projection method starts with an explicit estimate of the intermediate velocity field $u^{*}$
@@
where Re is the Reynolds number of the flow. Depending on the choice of integration method for the right-hand-side, we obtain explicit or semi-implicit methods. Here we will focus on fully explicit methods.
@@
The main contribution of this manuscript is injecting the Biot-Savart integral into the classical projection scheme and solving the resulting coupled problem. The momentum update using the new Biot-Savart domain condition proceeds as before: forming the intermediate velocity with Eq. 10 and projecting it to be divergence free with Eq. 12. However, the new boundary condition introduces a coupling between the body and domain boundaries. To see this, we substitute the update equation for the velocity at the new time-step (Eq. 12) into the Biot-Savart equation for the boundary velocity (Eq. 8) to give to complete equation for the boundary velocity at $t+\Delta t$
While the gradient of a field cannot induce curl in an unbounded fluid region (×$\nabla\times\nabla\phi\equiv 0$), the pressure generates a thin layer of vorticity on the body boundary $\partial\cal B$ every time-step [21]. This means that the last term in Eq. 14 is non-zero, introducing a non-local coupling between every point on the body boundary and every point on the domain boundary.
Because of the pressure contribution in Eq. 14, the pressure projection (Eq. 11) becomes
$\nabla\cdot\left[\frac{\Delta t}{\rho}\nabla p+f\left(\vec{x_{b}},\nabla\times\frac{\Delta t}{\rho}\nabla p,\Omega\right)\right]=\nabla\cdot\left[\vec{u}^{*}+f(\vec{x_{b}},\nabla\times\vec{u}^{*},\Omega)\right],\quad\forall\ \begin{array}[]{l}\vec{x}\in\Omega(N),\\ \vec{x}_{b}\in\partial\Omega(N^{\mathcal{S}})\end{array}$ (15)
The linear operator resulting from discretizing Eq. 11 with a local finite-difference finite-volume or finite-element method has only $O(N)$ non-zero terms, but the coupling induced by the Biot-Savart equation results in a dense matrix with at least $O(N^{1+S})$ non-zero terms. Note that this coupling is similar to the additional terms present in the Poisson equation when correctly imposing pressure boundary conditions for immersed boundary methods; see, for example, [22, 17]. Failing to correctly enforce the boundary conditions during projection leads to large errors in the predicted pressure and forces.
We must accurately include the coupling term in Eq. 14, but we want to avoid constructing the dense matrix or requiring substantial changes to existing Poisson solvers to make it practical for Eulerian schemes to use Biot-Savart boundary conditions. This is possible through a iterative operator-split method akin to the James-Lackner algorithm for potential flows
$\nabla\cdot\frac{\Delta t}{\rho}\nabla p^{k+1}=\nabla\cdot\left[\vec{u}^{*}+f\left(\vec{x_{b}},\nabla\times\left[\vec{u}^{*}-\frac{\Delta t}{\rho}\nabla p^{k}\right],\Omega\right)\right],\quad\forall\ \begin{array}[]{l}\vec{x}\in\Omega(N),\\ \vec{x}_{b}\in\partial\Omega(N^{\mathcal{S}}),\end{array}$ (16)
using the previous pressure solution $p^{k}$ to treat the Biot-Savart term as a residual and iterating until both the incompressible and boundary conditions are satisfied. Eulerian solvers typically use iterative methods to determine the pressure field in the projection step, making this a natural way to incorporate the coupling between the domain and body boundaries, without modifying the Poisson solver itself. In this work, we use a multigrid solver for the Poisson system, as described in [23], and updating the domain velocities and corresponding Poisson residuals at the end of each V-cycle is sufficient to ensure that the final pressure and velocity field without increasing the number of iterations.
Our integration of Eq. 16 into the unsteady momentum update significantly improves the stability of the coupled system compared to similar applications in potential flow such as  [7]. The error introduced by the operator partitioning is bounded by the change in pressure on the body over a single time step. We therefore never require relaxation to achieve convergence, even when impulsively starting a flow from rest with domain boundaries a small fraction of the body length away from the body boundaries. In addition, the resulting coupled system is remarkably robust to errors in the velocity reconstruction because the Biot-Savart equation is only used to set the boundary conditions, not update the entire flow field as in vortex-based methods. Consider the effect of neglecting the contribution from a vorticity source $\Delta\omega$ that has just left the domain, meaning this contribution is missing from ×$\nabla\times\vec{u}^{*}$ in Eqs.14 and 16. As only the boundary velocities are updated in Eq. 14, the momentum errors in the local Eulerian Navier-Stokes update are limited to the downwind characteristics from the domain boundary, and the impact on the (critical) inflow boundary regions are typically the smallest, as they are the farthest from the vorticity leaving the outflow boundary regions.